# Bumblebee gravity: spherically-symmetric solutions away from the potential minimum

### v1

@@ VI.1 Observational constraints
However, we discuss a general strategy could be used to measure parameters in this version of the bumblebee model, that involves orbital studies outside of the horizon where a post-Newtonian approach could be adopted. Generally speaking, the effects of black hole, apart from direct measurement of gravitational waves with black hole or other binary systems, are observed through its effects on orbiting bodies, including light. As an example, consider the precision observations of high speed stars around the central black hole of the Milky Way [123]. The post-Newtonian acceleration for an orbiting body at radius $r$ takes the form:
@@
where $\delta\vec{a}$ contains all relevant effects (e.g., other nearby bodies, relativistic effects, beyond GR effects, in a post-Newtonian series, etc. as in standard references like Ref.  [2]). Indeed, this type of equation has been used for orbital studies in a plethora of works to constrain local Lorentz violation in gravity before (e.g., [78, 41, 40, 124, 125]). Here we are interested in the form of perturbations $\delta\vec{a}$ stemming from the solutions found in this work in section V.
@@
Consider the case of the solution in Figure 6(d), the horizon expansion, for the hypergeometric potential $V=g\Lambda^{4}(M(4,2,X/\Lambda^{2})+1/e^{2})$ (which vanishes at the minimum value $X/\Lambda^{2}=-2$). This solution has a horizon, by construction, but also has two smoking gun features. One is the peculiar peak for $M$, and resulting “dip” in $N$, and the other is the steady decrease to zero afterwards, both in stark contrast to the standard behavior of the RN solution. Some preliminary work on plotting orbits shows the effect of such a metric. In Fig. 11, we show null geodesic trajectories of the RN metric and the $M(4,2,X/\Lambda^{2})$ metric of Fig. 6(d), with the same initial values of position and velocity. One can see the exaggerated features of the orbit in the symmetry-breaking case (a hypergeometric potential). We leave a thorough analysis of orbits for future work.
@@
Figure 11: Sample null geodesic trajectories of the RN metric and the $M(4,2,X/\Lambda^{2})$ metric of Fig. 6(d). The black circle is the horizon and the axes are the spherical projections onto the orbital plane.
@@
Following up on this example, one could crudely estimate $\delta\vec{a}$ from examining the weak field geodesic equation for the radial acceleration, $a^{r}\approx-\Gamma^{r}_{\phantom{r}tt}\approx\tfrac{1}{2}\partial_{r}g_{tt}$. In terms of $N(x)$, this would be $a^{r}=-(1/2r_{h})dN/dx$. Measurements of accelerations of objects around stars can yield results for orbital quantities indicating errors at the $1\%$ level or less [126]. Thus, we could expect to place limits on acceleration perturbations at this level. What is needed, for the horizon expansion solutions, is a numerical study of the dependence of $dN/dx$ on the values of three relevant parameters $g$ and $k$, and the initial value $a_{f1}$. One could then use data to extract a statistical distribution for these parameters.
@@
Some study of the variation of the three parameters for this case shows that for $k\approx 3\times 10^{-5}$, $a_{f1}=1/10$, and ${\tilde{g}}=1/100$ results in a $\Delta a/a\sim 1/100$ where $\Delta a$ is the difference in the radial accelerations between the RN solution and the solution with the $M(4,2,X/\Lambda^{2})$ potential case. Note that these are not remarkably small values, as is typically suggested for spacetime symmetry breaking [127]. Indeed, previous studies have shown that large Lorentz violation is indeed possible in some sectors even with present experimental limits [128, 28, 98, 129].
@@
For the asymptotic expansion approach of Section 70, similar approaches as described for the horizon expansion could be adopted. That is, one explores the effective acceleration of a test body numerically, as a function of the parameters ${\tilde{g}}$, $k$, and in this case $c_{f0}$ and $c_{f1}$.
@@ VI.2 Summary
Also in flat spacetime, we discussed the solution to the bumblebee field equation in the spherical symmetry case in Section III.2.2. The perturbative solution around the minimum $b_{\mu}$ was obtained in (20). For the fully nonlinear case, we characterized the asymptotic behavior of the solutions for different choices of initial conditions in (30)-(32), which can be seen also in Fig. 2-4.
In Section IV, we investigated some general aspects of the field equations with gravity. First, we briefly discussed energy conditions, followed by a sample solution showing that AdS spacetime could arise in some cases in (42). Then, in Sec. IV.3 we examined the Bekenstein horizon argument, resulting in (53), and concluded there are no constraints on the bumblebee model with arbitrary potentials. Another important set of results are the special case solutions, where analytical solutions are possible in Sec. IV.4. The chart in Fig. 5 categorized these solutions.
Numerical solutions to the full equations were the subject of Sec. V. Firstly, we established the field equations in Eddington-Finkelstein coordinates in (65) and (66). In Sec. V.1, we then focused on the horizon region and the asymptotic regions, using two different series expansions (68) and (70), with lower-order coefficients in Tables 2 and 3. These expansions were then used to “seed” a full numerical solver. Sample solutions for different choices of boundary values are shown in Fig. 6 for the horizon expansion and Fig. 7 for the asymptotic expansion.
We primarily studied four types of potentials: Proca (massive vector), quadratic, and two hypergeometric choices $M(n,2,X/\Lambda^{2})$ for $n=3$ and $n=4$. General features observed include sensitivity to initial values at the horizon $r=r_{h}$ or at infinity $x\rightarrow\infty$. We found naked singularities as possibilities, as illustrated in Fig. 7(b). Also we found wildly varying behavior of the metric functions (equivalent to changing signs of $\partial_{r}g_{tt}$ some distance from the horizon), as illustrated in Fig. 6(d). In addition, the curvature evaluated at the horizons or at infinity, was found to diverge, or converge, depending on the type of potential, and the initial values chosen. The latter can be seen in Figs. 8(a) and 8(b).
@@
We view this work as a “door opening” into the rich structure of possibilities of solutions with $V^{\prime}\neq 0$. Since we showed that such models with hypergeometric potential functions do not automatically result in energies unbounded from below, it represents a complimentary area to explore, alongside literature on the couplings $\xi_{1}$ and $\xi_{2}$. It would be of interest to study these types of solutions further. As already outlined above, constructing a numerical scheme that varies the parameters in the model, while calculating an observable, such as an orbital effect, and then comparing against orbital data would be of interest. This may involve a step-up in computing time, given that the numerical solver used in this work was at least mildly time-consuming.
@@ Acknowledgements.
The authors thank Albin Nilsson and Robertus Potting for useful discussions that improved the manuscript. For Q.G.B., H.S.M. and D.W-C. this work was supported by the National Science Foundation grant number 2308602. H.S.M. and D.W-C. were supported by the Embry-Riddle Aeronautical University Undergraduate Research Institute and the Arizona NASA Space Grant consortium.

### v2

@@ VI.1 Observational constraints
However, we discuss a general strategy that could be used to measure parameters in this version of the bumblebee model, that involves orbital studies outside of the horizon where a post-Newtonian approach could be adopted. Generally speaking, the effects of a black hole, apart from direct measurements of gravitational waves from black hole binary systems, are observed through its effects on orbiting bodies, including light. As an example, consider the precision observations of high speed stars around the central black hole of the Milky Way [123]. The post-Newtonian acceleration for an orbiting body at radius $r$ takes the form:
@@
where $\delta\vec{a}$ contains all relevant effects (e.g., other nearby bodies, relativistic effects, beyond GR effects, in a post-Newtonian series, etc. as in standard references like Ref.  [2]). Indeed, this type of equation has been used for orbital studies in a plethora of works to constrain local Lorentz violation in gravity before (e.g., Refs. [78, 41, 40, 124, 125]). Here we are interested in the form of perturbations $\delta\vec{a}$ stemming from the solutions found in this work in Section V.
@@
Consider the case of the solution in Fig. 6d, the horizon expansion, for the hypergeometric potential $V=g\Lambda^{4}(M(4,2,z)+1/3e^{2})$ (which vanishes at the minimum value $z=X/\Lambda^{2}=-2$). This solution has a horizon, by construction, but also has two smoking gun features. One is the peculiar peak for $M$, and resulting “dip” in $N$, and the other is the steady decrease to zero afterwards, both in stark contrast to the standard behavior of the RN solution. Some preliminary work on plotting orbits shows the effect of such a metric. In Fig. 11, we show null geodesic trajectories of the RN metric and the $M(4,2,z)$ metric of Fig. 6d, with the same initial values of position and velocity. One can see the exaggerated features of the orbit in the symmetry-breaking case (a hypergeometric potential). We leave a thorough analysis of orbits for future work.
@@
Figure 11: Sample null geodesic trajectories of the RN metric and the $M(4,2,z)$ metric of Fig. 6d. The black circle is the horizon and the axes are the spherical projections onto the orbital plane with $\theta=\pi/2$.
@@
Following up on this example, one could crudely estimate $\delta\vec{a}$ from examining the weak field geodesic equation for the radial acceleration, $a^{r}\approx-\Gamma^{r}_{\phantom{r}tt}\approx\tfrac{1}{2}\partial_{r}g_{tt}$. In terms of $N(x)$, this would be $a^{r}=-(1/2r_{h})dN/dx$. Measurements of accelerations of objects around black holes can yield results for orbital quantities indicating errors at the $1\%$ level or less [126]. Thus, we could expect to place limits on acceleration perturbations at this level. What is needed, for the horizon expansion solutions, is a numerical study of the dependence of $dN/dx$ on the values of three relevant parameters $g$ and $k$, and the initial value $a_{f1}$. One could then use data to extract a statistical distribution for these parameters.
@@
Some study of the variation of the three parameters for this case shows that for $k\approx 3\times 10^{-5}$, $a_{f1}=1/10$, and ${\tilde{g}}=1/100$ results in a $\Delta a/a\sim 1/100$ where $\Delta a$ is the difference in the radial accelerations between the RN solution and the solution with the $M(4,2,z)$ potential case. Note that these are not remarkably small values, as is typically suggested for spacetime symmetry breaking [127]. Indeed, previous studies have shown that large Lorentz violation is possible in some sectors even with present experimental limits [128, 28, 99, 129].
@@
For the asymptotic expansion approach of Section V.1.2, similar approaches as described for the horizon expansion could be adopted. That is, one explores the effective acceleration of a test body numerically, as a function of the parameters $\tilde{g}$}, $k$, and in this case $c_{f0}$ and $c_{f1}$.
@@ VI.2 Summary
Also in flat spacetime, we discussed the solution to the bumblebee field equation in the spherical symmetry case in Section III.2.2. The perturbative solution around the minimum $b_{\mu}$ was obtained in equation (20). For the fully nonlinear case, we characterized the asymptotic behavior of the solutions for different choices of initial conditions in equations (30)-(32), which can be seen also in Figs. 2-4.
In Section IV, we investigated some general aspects of the field equations with gravity. First, we briefly discussed energy conditions, followed by a sample solution showing that AdS spacetime could arise in some cases in (42). Then, in Sec. IV.3 we examined the Bekenstein horizon argument, resulting in equation (53), and concluded there are no constraints on the bumblebee model with arbitrary potentials. Another important set of results are the special case solutions, where analytical solutions are possible in Sec. IV.4. The chart in Fig. 5 categorized these solutions.
Numerical solutions to the full equations were the subject of Sec. V. Firstly, we established the field equations in Eddington-Finkelstein coordinates in equations (65) and (66). In Sec. V.1, we then focused on the horizon region and the asymptotic regions, using two different series expansions (68) and (70), with lower-order coefficients in Tables 2-5. These expansions were then used to seed a full numerical solver. Sample solutions for different choices of boundary values are shown in Fig. 6 for the horizon expansion in Fig. 7 for the asymptotic expansion.
We primarily studied four types of potentials: Proca (massive vector), quadratic, and two hypergeometric choices $M(n,2,z)$ for $n=3$ and $n=4$. General features observed include sensitivity to initial values at the horizon $r=r_{h}$ or at infinity $x\rightarrow\infty$. We found naked singularities as possibilities, as illustrated in Fig. 7b. Also we found wildly varying behavior of the metric functions (equivalent to changing signs of $\partial_{r}g_{tt}$ some distance from the horizon), as illustrated in Fig. 6d. In addition, the curvature evaluated at the horizons or at infinity, was found to diverge, or converge, depending on the type of potential, and the initial values chosen. The latter can be seen in Figs. 8a and 8b.
@@
We view this work as a “door opening” into the rich structure of possibilities of solutions with $V^{\prime}\neq 0$. Since we showed that such models with hypergeometric potential functions do not automatically result in energies unbounded from below, it represents a complimentary area to explore, alongside literature on the couplings $\xi_{1}$ and $\xi_{2}$. It would be of interest to study these types of solutions further. As already outlined above, constructing a numerical scheme that varies the parameters in the model, while calculating an observable, such as an orbital effect, and then comparing against orbital data would be of interest. This may involve a step up in computing time, given that the numerical solver used in this work was at least mildly time-consuming on an ordinary computer.
@@ Acknowledgements.
The authors thank Nils A. Nilsson, Robertus Potting, and Edward Poon for useful discussions that improved the manuscript. For Q.G.B., H.S.M. and D.W.-C. this work was supported by the National Science Foundation grant number 2308602. H.S.M. and D.W.-C. were supported by the Embry-Riddle Aeronautical University Undergraduate Research Institute and the Arizona NASA Space Grant consortium.